use the given conditions to write and equation for each line in point-slope form and slope intercept form
passing through(3,5) and (8,15)

Answer by lwsshak3(11628)

use the given conditions to write and equation for each line in point-slope form and slope intercept form
passing through(3,5) and (8,15)
**
slope,m=∆y/∆x=(y1-y2)/x1-x2)=(15-5)/(8-3)=10/5=2
point slope form:
(y1-y2)=m(x1-x2)
Using one of the given points,
y-5=2(x-3)
y-5=2x-6
..
slope intercept form:
y=mx+b
y=2x+b
solving for b using one of the given points,
5=2(3)+b
b=-1
y=2x-1